OpenStudy (anonymous):

the number pi^3-4pi is equal to?

OpenStudy (anonymous):

If that is officially pi^(3-4*pi) then the answer is 1.754203879*10^-5. Is that enough to answer your question?

OpenStudy (anonymous):

no. the choices that were given to me were (pi-2)pi(pi+2), pi^-1 and -pi

OpenStudy (anonymous):

oh and the question is suppose to be written (pi^3)-4pi. if that's clearer.

OpenStudy (anonymous):

There's an infinite number of things that equal to (pi^3)-4pi. Next time specify right away what the question is precisely asking and the possible answers, if given (pi^3)-4pi = (pi^2 - 2pi)(pi+2) (pi^2 - 2pi)(pi+2) = (pi-2)pi(pi+2) So first answer correct